# Some Constructions on Quantum Principal Bundles

### v1

@@ Abstract.
This paper aims to prove four statements in the theory of quantum principal bundles:
@@
    The universal differential envelope $\ast$–calculus of a matrix (compact) Lie group, for the classical bicovariant $\ast$–First Order Differential Calculus, is the algebra of differential forms.
@@ 1. Universal Differential Envelope $\ast$–calculus in Differential Geometry
Let $(\Gamma,d)$ be a bicovariant $\ast$–First Order Differential Calculus ($\ast$–FODC) over $\mathcal{G}$ ([3, 6]) and consider the $\mathbb{C}$–vector space given by
@@
with $X^{2}=\{\displaystyle\sum^{n}_{i=1}a_{i}\,b_{i}\mid a_{i},\,b_{i}\,\in\,X\,\mbox{ for some }\,n\,\in\,\mathbb{N}\}$. Moreover, for every $f$in $X$,
@@
where $[f]_{X^{2}}$ denotes the equivalence class of $f$ in $X/X^{2}$. In this way, for every $g$in $\mathrm{Ker}(\epsilon)$ we get $dg_{e}=[g]_{\mathcal{R}}=[g]_{X^{2}}$. By this relation, $g$in $\mathcal{R}$ if and only if $g$in $=X^{2}\cap\mathrm{Ker}(\epsilon)=\mathrm{Ker}^{2}(\epsilon),$ which proves that $\mathcal{R}=\mathrm{Ker}^{2}(\epsilon)$.
@@ 4. The form of the maps $\{T^{\mathrm{L}}_{k}\}$
Let $G$ be a compact matrix Lie group and let $\pi:P\longrightarrow B$ be a classical principal $G$–bundle, where $P$ is the total space, $B$ is the base space and $\pi$ is the bundle projection. Assume $P$, $B$ are compact. If $\mathcal{T}$ is a complete set of mutually non–equivalent irreducible $\mathcal{G}$–corepresentations with $\delta^{\mathbb{C}}_{\mathrm{triv}}$in $\mathcal{T}$, then for every $\delta^{V}$in $\mathcal{T}$, there exists $\{T^{\mathrm{L}}_{k}\}^{d_{V}}_{k=1}$subseteq $\textsc{Mor}(\delta^{V},\gamma^{\#})$, for some $d_{V}$in $\mathbb{N}$, such that equation (30) holds. Here
$\gamma^{\#}:C^{\infty}_{\mathbb{C}}(P)\longrightarrow C^{\infty}_{\mathbb{C}}(P\times G)\cong C^{\infty}_{\mathbb{C}}(P)\otimes C^{\infty}_{\mathbb{C}}(G)$
@@
Let $\delta^{V}$in $\mathcal{T}$ and consider its associated representation $\alpha^{V}$. Since $\alpha^{V}$ is unitary, the associated vector bundle $\pi_{\alpha^{V}}:E^{V}\longrightarrow B$ is a Hermitian bundle.
For each $b$in $B$, let $(U_{b},\Phi_{b})$ be a local trivialization of the Hermitian bundle $\pi_{\alpha^{V}}:E^{V}\longrightarrow B$ around $b$ associated to a principal $G$–bundle local trivialization $(U_{b},\Psi_{b})$ of $\pi:P\longrightarrow B$. Then there exists a set of local sections $\{\hat{s}^{b}_{1},\ldots,\hat{s}^{b}_{1}\}$subseteq $\Gamma\bigl{(}\pi_{{\alpha^{V}}}^{-1}(U_{b})\bigr{)}$ such that $\hat{s}^{b}_{i}(a)=(a,e_{i})$ for all $a$in $U_{b}$ (under the diffeomorphism $\Phi_{b}$). Since $\{U_{b}\}_{b\in B}$ is an open cover, by compactness, there exist points $b_{1},\ldots,b_{r}\in B$ such that $\{\,U_{b_{i}}\}_{i=1}^{r}$ remains an open cover of $B$. Let $\{\rho_{b_{i}}\}_{i=1}^{r}$ be a partition of unity with $\rho_{b_{i}}$ having compact support and subordinate to the open cover $\{\,U_{b_{i}}\}_{i=1}^{r}$. In this way, consider the global sections of $E^{V}$
@@
For every $x$in $P$, if $a=\pi(x)$in $B$, it is well–known that there exists a unique $v$in $V$ such that $s^{i}_{j}(a)=[x,v]$in $E^{V}$. This induces a smooth map
$f^{i}_{j}:P$ $\longrightarrow V$
$x$ $\longrightarrow v=\sqrt{\rho_{b_{i}}(\pi(x))}\;\alpha^{V}(A(x)^{-1},e_{j}),$
where $A(x)$ is the unique element of $G$ such that $\Psi_{b_{i}}(\pi(x),A(x))=x$. These maps satisfy $f^{i}_{j}(\gamma(x,B))=\alpha^{V}(B^{-1},f^{i}_{j}(x))$ for every $x$in $P$, $B$in $G$. We define the smooth functions
@@
$\left(\sum^{r,n}_{l,k=1}f^{l\,\ast}_{ki}f^{l}_{kj}\right)(x)=\delta_{ij},\qquad(\gamma^{\#}\circ T^{i}_{j})(x,B)=((T^{i}_{j}\otimes\mathrm{id}_{H})\circ\delta^{V})(x,B).$

### v2

@@ Abstract.
This paper works as an appendix of the paper titled Geometry of Associated Quantum Vector Bundles and the Quantum Gauge Group. Here, we are going to prove four statements in the theory of quantum principal bundles:
@@
    The universal differential envelope $\ast$–calculus of a matrix (compact) Lie group, for the classical bicovariant First Order Differential $\ast$–Calculus, is the algebra of differential forms.
@@ 1. Universal Differential Envelope $\ast$–calculus in Differential Geometry
Let $(\Gamma,d)$ be a bicovariant First Order Differential $\ast$–Calculus ($\ast$–FODC) over $\mathcal{G}$ ([3, 6]) and consider the $\mathbb{C}$–vector space given by
@@
with $X^{2}=\{\displaystyle\sum^{n}_{i=1}a_{i}\,b_{i}\mid a_{i},\,b_{i}\,\in\,X\,\mbox{ for some }\,n\,\in\,\mathbb{N}\}$ [7]. Moreover, for every $f$in $X$,
@@
where $[f]_{X^{2}}$ denotes the equivalence class of $f$ in $X/X^{2}$ [7]. In this way, for every $g$in $\mathrm{Ker}(\epsilon)$ we get $dg_{e}=[g]_{\mathcal{R}}=[g]_{X^{2}}$. By this relation, $g$in $\mathcal{R}$ if and only if $g$in $=X^{2}\cap\mathrm{Ker}(\epsilon)=\mathrm{Ker}^{2}(\epsilon),$ which proves that $\mathcal{R}=\mathrm{Ker}^{2}(\epsilon)$.
@@ 4. The form of the maps $\{T^{\mathrm{L}}_{k}\}$
Let $G$ be a compact matrix Lie group and let $\pi:P\longrightarrow B$ be a classical principal $G$–bundle, where $P$ is the total space, $B$ is the base space and $\pi$ is the bundle projection. Assume $P$, $B$ are compact. If $\mathcal{T}$ is a complete set of mutually non–equivalent irreducible $\mathcal{G}$–corepresentations with $\delta^{\mathbb{C}}_{\mathrm{triv}}$in $\mathcal{T}$, then for every $\delta^{V}$in $\mathcal{T}$, there exists $\{T^{\mathrm{L}}_{k}\}^{d_{V}}_{k=1}$subseteq $\textsc{Mor}(\delta^{V},\Delta_{P})$, for some $d_{V}$in $\mathbb{N}$, such that equation (30) holds. Here
$\Delta_{P}:C^{\infty}_{\mathbb{C}}(P)\longrightarrow C^{\infty}_{\mathbb{C}}(P\times G)\cong C^{\infty}_{\mathbb{C}}(P)\otimes C^{\infty}_{\mathbb{C}}(G)$
@@
Let $\delta^{V}$in $\mathcal{T}$ and consider its associated representation $\alpha^{V}$ with $n=\mathrm{dim}_{\mathbb{C}}(V)$. Since $\alpha^{V}$ is unitary, the associated vector bundle $\pi_{\alpha^{V}}:E^{V}\longrightarrow B$ is a Hermitian bundle.
For each $b$in $B$, let $(U_{b},\Phi_{b})$ be a local trivialization of the Hermitian bundle $\pi_{\alpha^{V}}:E^{V}\longrightarrow B$ around $b$ associated to a principal $G$–bundle local trivialization $(U_{b},\Psi_{b})$ of $\pi:P\longrightarrow B$. Then there exists a set of local sections $\{\hat{s}^{b}_{1},\ldots,\hat{s}^{b}_{1}\}$subseteq $\Gamma\bigl{(}\pi_{{\alpha^{V}}}^{-1}(U_{b})\bigr{)}$ such that $\hat{s}^{b}_{i}(a)=(a,e_{i})$ for all $a$in $U_{b}$ (under the diffeomorphism $\Phi_{b}$). Since $\{U_{b}\}_{b\in B}$ is an open cover, by compactness, there exist points $b_{1},\ldots,b_{r}\in B$ such that $\{\,U_{b_{i}}\}_{i=1}^{r}$ remains an open cover of $B$. Let $\{\rho_{b_{i}}\}_{i=1}^{r}$ be a partition of unity subordinate to the open cover $\{\,U_{b_{i}}\}_{i=1}^{r}$, where each $\rho_{b_{i}}$ has compact support and admits a smooth square root. In this way, consider the global sections of $E^{V}$
@@
Notice that $\{s^{i}_{j}\}^{r,n}_{i,j=1}$ is a set of $\mathbb{C}^{\infty}_{\mathbb{C}}(B)$–bimodule generators of $\Gamma(E^{V})$.
It is well–known that $\Gamma(E^{V})$ is isomorphic to the space of $G$–equivariant smooth functions $C^{\infty}_{\mathbb{C}}(P,V)^{G}=\{f:P\longrightarrow V\mid f\mbox{ is smooth such that }f(xA)=\alpha(A^{-1})f(x)\}$ as $\mathbb{C}^{\infty}_{\mathbb{C}}(B)$–bimodules. Then the maps
$f^{i}_{j}:P$ $\longrightarrow V,$
$x$ $\longrightarrow v=\sqrt{\rho_{b_{i}}(\pi(x))}\;\alpha(A(x)^{-1},e_{j}),$
where $A(x)$ is the unique element of $G$ such that $\Psi_{b_{i}}(\pi(x),A(x))=x$, from a set of $\mathbb{C}^{\infty}_{\mathbb{C}}(B)$–bimodule generators of $C^{\infty}_{\mathbb{C}}(P,V)^{G}$ associated to $\{s^{i}_{j}\}^{r,n}_{i,j=1}$. We define the smooth functions
@@
$\left(\sum^{r,n}_{l,k=1}f^{l\,\ast}_{ki}f^{l}_{kj}\right)(x)=\delta_{ij},\qquad(\Delta_{P}\circ T^{i}_{j})(x,A)=((T^{i}_{j}\otimes\mathrm{id}_{H})\circ\delta^{V})(x,A).$
@@
It is worth mentioning that the maps $T^{i}_{j}$ agree with the dualization of $f^{i}_{j}$ via the pull–back, once the dual space $V^{\#}$ is identified with $V$. Moreover, to define the maps $\{T^{\mathrm{L}}_{i}\}^{d_{V}}_{i=1}$, it was necessary to have a partition of unity on the base space with smooth square roots. This is why having a $C^{\ast}$–algebra as the quantum base space (or one that can be completed to a $C^{\ast}$–algebra) is a sufficient condition to guarantee the existence of the maps $\{T^{\mathrm{L}}_{i}\}_{i=1}^{d_{V}}$ in the non–commutative geometrical setting, as the reader can verify in [8].